# Exact Bayesian Inference for Markov Switching Diffusions

### v1

@@ 7 Discussion
In spite of those advances, we have restricted our discussion and experiments to univariate diffusions, even as applications often call for multivariate models. While the methodology may be in principle extended to cover some of those models, the conditions that allow for a Lamperti transform are more stringent in multivariate models, as pointed out in beskos2008factorisation. Moreover, the Brownian bridge proposals considered herein work best on diffusions for which the Lamperti-transformed process $X$ has state space $\mathcal{X}=\mathbf{R}$. Covering those cases would further exacerbate the already significant computational and methodological commitment that is required in the more restrictive setting of this paper.
In terms of more immediate avenues of research, we note that $\lambda$ can be constructed to give rise to various behaviors - for instance, such that $Y$ approximates a continuous-state process. This is naturally accommodated by our methods, and we are currently exploring some of those variations. Another possibility is to address the case of semi-Markov switching diffusions, where the latent process has non-exponential holding times. Most of our algorithm could be applied out the box, though modifications would have to be made to the localized hidden data update described in Supplement B, since bridge simulation of the latent process is more complicated, and different sections of the trajectory cannot be updated independently.
@@ Acknowledgements
The authors are grateful to Flávio Gonçalves for his comments on the manuscript.
Timothée Stumpf-Fétizon has received funding from the European Research Council (ERC, PrSc-HDBayLe, Grant agreement No. 101076564). Krzysztof Łatuszynski has been supported by the Royal Society through the Royal Society University Research Fellowship. Gareth O. Roberts has been supported by the UKRI grant EP/Y014650/1 as part of the ERC Synergy project OCEAN, EPSRC grants Bayes for Health (R018561), CoSInES (R034710), PINCODE (EP/X028119/1), ProbAI (EP/Y028783/1) and EP/V009478/1.
[…]
@@ Appendix A Proof of Proposition 1
$\mathop{}\!\mathrm{d}X_{t}=\delta_{\theta}(X_{t},y_{\dot{\tau}})\mathop{}\!\mathrm{d}t+\rho_{\theta}(y_{\dot{\tau}})\mathop{}\!\mathrm{d}W_{t},\qquad(X_{0}=\eta_{\theta}(v_{0}),\quad t\in[\dot{\tau},\ddot{\tau}))$ (64)
$\delta_{\theta}(a,b)=\left\lparen\frac{\mu_{\theta}(\cdot,b)}{\sigma_{\theta}}-\frac{\sigma_{\theta}^{\prime}}{2}\right\rparen\circ\eta_{\theta}^{-1}(a).$ ) . (65)
@@
$\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau}]})$ $=\exp\left\{\int_{\dot{\tau}}^{\ddot{\tau}}\frac{\delta_{\theta}(x_{t},y_{\dot{\tau}})}{\rho_{\theta}(y_{\dot{\tau}})}\mathop{}\!\mathrm{d}W_{t}+\frac{1}{2}\int_{\dot{\tau}}^{\ddot{\tau}}\frac{\delta_{\theta}^{2}(x_{t},y_{\dot{\tau}})}{\rho_{\theta}^{2}(y_{\dot{\tau}})}\mathop{}\!\mathrm{d}t\right\}.$ (66)
@@
$\Delta_{\theta}(a,b)=\int_{v^{*}}^{a}\delta_{\theta}(c,b)\mathop{}\!\mathrm{d}c,\qquad(v^{*}\in\mathcal{V})$ (67)
@@
$\varphi_{\theta}(a,b)=\frac{1}{2}\left\lparen\frac{\delta_{\theta}^{2}(a,b)}{\rho_{\theta}^{2}(b)}+\partial_{a}\delta_{\theta}(a,b)\right\rparen.$ ) . (70)
@@
$\begin{split}\pi(x_{\ddot{\tau}}|x_{\dot{\tau}},y_{\dot{\tau}},\theta)\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(x_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(x_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau})})&=\operatorname{\mathrm{N}}\left[x_{\ddot{\tau}};x_{\dot{\tau}},(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau}]}).\end{split}$ (73)
@@
and changing variables back to $V_{\ddot{\tau}}$, the final expression is
@@
×$\qquad\times\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau})},\eta_{\theta}(v_{\ddot{\tau}})).$
@@
$\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\rho_{\theta}(y_{\dot{\tau}})z_{t}+\eta_{\theta}(v_{\dot{\tau}})+(\eta_{\theta}(v_{\ddot{\tau}})-\eta_{\theta}(v_{\dot{\tau}}))\frac{t-\dot{\tau}}{\ddot{\tau}-\dot{\tau}}\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (77)
@@
$\begin{split}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)&=|\eta_{\theta}^{\prime}(v_{\ddot{\tau}})|\operatorname{\mathrm{N}}\left[\eta_{\theta}(v_{\ddot{\tau}});\eta_{\theta}(v_{\dot{\tau}}),(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]\\ &\qquad\times\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}(\zeta_{\theta}^{-1}(z_{(\dot{\tau},\ddot{\tau})};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),\eta_{\theta}(v_{\ddot{\tau}}))\\ &=|\eta_{\theta}^{\prime}(v_{\ddot{\tau}})|\operatorname{\mathrm{N}}\left[\eta_{\theta}(v_{\ddot{\tau}});\eta_{\theta}(v_{\dot{\tau}}),(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]e^{\Delta_{\theta}(\eta_{\theta}(v_{\ddot{\tau}}),y_{\dot{\tau}})-\Delta_{\theta}(\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}})}\\ &\qquad\times\exp\left\{-\int_{\dot{\tau}}^{\ddot{\tau}}\varphi_{\theta}(\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),y_{\dot{\tau}})\mathop{}\!\mathrm{d}t\right\},\end{split}$ (79)
@@ Appendix B Localizing and Adapting the Hidden Data Update
The approach consists of conditioning the $(V_{R},Z,Y)$-update on $Y_{N}$, where $N\subseteq(s\mathbin{\backslash}\{0,\omega\})$ is a random set of times. If no element of $s$ is included in $N$ with probability 1, this update may be thought of as a random scan Gibbs update, which preserves ergodicity of the Markov chain. The main rationale to limiting conditioning times to elements of $s$ is that in its EA2/EA3 representation, $Z$ is only semi-Markovian at times $\tau$, and therefore the full conditional does not factorize neatly at times $\nu\not\subset\tau$.
With $\nu$ fixed, we generate updates according to $\pi(v_{r},y\mathbin{\backslash}y_{\nu}|v_{s},y_{\nu},\theta,\lambda)$. On the one hand, the conditional proposal $\kappa(v_{r},y\mathbin{\backslash}y_{\nu}|v_{s},y_{\nu})$ has a smaller step size, thereby increasing the acceptance probability. On the other hand, by the Markov property, we immediately benefit from the factorization
@@
with an analogous factorization for $\pi(v_{r},z,y\mathbin{\backslash}y_{\nu}|v_{s},y_{\nu},\theta,\lambda)$, so generation and acceptance of the proposal may be partitioned according to $\nu$. This further increases the acceptance probability, and reduces Bernoulli factory run time. The proposal law
@@
$\begin{split}&\kappa(v_{r^{\dagger}\cap(0,\nu_{1})}^{\dagger},z_{(0,\nu_{1})}^{\dagger},y_{[0,\nu_{1})}^{\dagger}|v_{s},y_{\nu_{1}})\\ &\quad=\pi(y_{[0,\nu_{1})}^{\dagger}|y_{\nu_{1}},\lambda)\kappa(v_{r^{\dagger}\cap(0,\nu_{1})}^{\dagger}|v_{s},y^{\dagger})\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}\cap[0,\nu_{1}]}\kappa(z_{(\dot{\tau},\ddot{\tau})}^{\dagger}),\\ \end{split}$ (83)
@@
merely involve the backward and forward law $\pi(y_{[0,\nu_{1})}^{\dagger}|y_{\nu_{1}},\lambda)$ and $\pi(y_{(\nu_{|\nu|},\omega]}^{\dagger}|y_{\nu_{|\nu|}},\lambda)$ respectively. Simulation of the Markov jump process bridges from $\pi(y_{(\dot{\nu},\ddot{\nu})}^{\dagger}|y_{\{\dot{\nu},\ddot{\nu}\}},\lambda)$ may be carried out according to any of the schemes proposed in hobolth2009simulation. We further observe the decomposition
@@
$\begin{split}\frac{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\},\{v_{r},z,y\})}{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\{v_{r},z,y\},\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\})}&=\prod_{(\dot{\varsigma}\sim\ddot{\varsigma})\in\varsigma\cap\{0,\omega\}}\frac{\kappa(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}|v_{s},y_{\varsigma})}{\kappa(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger}|v_{s},y_{\varsigma})}\frac{\pi(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger}|v_{s},y_{\varsigma},\theta,\lambda)}{\pi(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger}|v_{s},y_{\varsigma},\theta,\lambda)}\\ &=\prod_{(\dot{\varsigma}\sim\ddot{\varsigma})\in\varsigma\cap\{0,\omega\}}\frac{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger},\gamma_{(\dot{\varsigma},\ddot{\varsigma})})}{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\gamma_{(\dot{\varsigma},\ddot{\varsigma})},\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger})},\end{split}$ (87)

### v2

@@ 7 Discussion
In spite of those advances, we have restricted our discussion and experiments to univariate diffusions, even as applications often call for multivariate models. While the methodology may be in principle extended to cover some of those models, the conditions that allow for a Lamperti transform are more stringent in multivariate models, as pointed out in [5]. Moreover, the Brownian bridge proposals considered herein work best on diffusions for which the Lamperti-transformed process $X$ has state space $\mathcal{X}=\mathbf{R}$. Covering those cases would further exacerbate the already significant computational and methodological commitment that is required in the more restrictive setting of this paper.
Furthermore, the methodological breadth and generality of our algorithm comes at a cost in theoretical tractability, which we consider beyond the scope of this paper. Since many of the methods involved do not have uniform characteristics over unbounded parameter spaces, we believe that an analysis of MCMC convergence rates would require strong restrictions on the model class or the prior. Nonetheless, there is theoretical support for some components of the algorithm, such as for Gibbs sampling of $(Z,\Theta)$ in ordinary Ito diffusions, for which [40] found that irreducibility of the Gibbs sampler is synonymous with “good mixing”. Furthermore, for fixed $\theta$, modified versions of our independence sampler for the hidden data can be shown to be uniformly ergodic for various model classes through an argument along the lines of [35]. This usually requires a more heavy-tailed proposal for $V_{R}$ (diffusion value at jump times) to dominate unbounded terms in Equation (9), for instance from a Cauchy process. In particular, uniform ergodicity can be shown for models with bounded drift, such as the Tanh diffusion in Section 5, and Ornstein-Uhlenbeck type models with up to linear drift. Conversely, models with exponential drift such as the switching logistic growth diffusion from Section 6 pose greater challenges.
As for more immediate avenues of research, we note that $\lambda$ can be constructed to give rise to various behaviors - for example, such that $Y$ approximates a continuous-state process [29, 13]. This is naturally accommodated by our methods, and we are currently exploring some of those variants. Another possibility is to address the case of semi-Markov switching diffusions, where the latent process has non-exponential holding times. Most of our algorithm could be applied out the box, though modifications would have to be made to the localized hidden data update described in Supplement B, since bridge simulation of the regime process is more complicated, and different sections of the trajectory cannot be updated independently.
@@ Acknowledgements
The authors are grateful to Flávio Gonçalves for his comments on the manuscript. Timothée Stumpf-Fétizon has received funding from the ERC grant PrSc-HDBayLe (No. 101076564). Krzysztof Łatuszynski has been supported by the Royal Society through the Royal Society University Research Fellowship. Gareth O. Roberts has been supported by the UKRI grant OCEAN (EP/Y014650/1) and EPSRC grants Bayes for Health (R018561), CoSInES (R034710), PINCODE (EP/X028119/1), ProbAI (EP/Y028783/1) and EP/V009478/1.
@@ Appendix A Proof of Proposition 1
$\mathop{}\!\mathrm{d}X_{t}=\delta_{\theta}(X_{t},y_{\dot{\tau}})\mathop{}\!\mathrm{d}t+\rho_{\theta}(y_{\dot{\tau}})\mathop{}\!\mathrm{d}W_{t},\qquad(X_{0}=\eta_{\theta}(v_{0}),\;t\in[\dot{\tau},\ddot{\tau}))$ (62)
[…]
@@
$\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau}]})$ $=\exp\left\{\int_{\dot{\tau}}^{\ddot{\tau}}\frac{\delta_{\theta}(x_{t},y_{\dot{\tau}})}{\rho_{\theta}(y_{\dot{\tau}})}\mathop{}\!\mathrm{d}W_{t}+\frac{1}{2}\int_{\dot{\tau}}^{\ddot{\tau}}\frac{\delta_{\theta}^{2}(x_{t},y_{\dot{\tau}})}{\rho_{\theta}^{2}(y_{\dot{\tau}})}\mathop{}\!\mathrm{d}t\right\}$ (64)
$=\exp\left\{\int_{\dot{\tau}}^{\ddot{\tau}}\frac{\delta_{\theta}(x_{t},y_{\dot{\tau}})}{\rho_{\theta}^{2}(y_{\dot{\tau}})}\mathop{}\!\mathrm{d}X_{t}-\frac{1}{2}\int_{\dot{\tau}}^{\ddot{\tau}}\frac{\delta_{\theta}^{2}(x_{t},y_{\dot{\tau}})}{\rho_{\theta}^{2}(y_{\dot{\tau}})}\mathop{}\!\mathrm{d}t\right\}.$
@@
$\Delta_{\theta}(a,b)=\int_{x^{*}}^{a}\delta_{\theta}(c,b)\mathop{}\!\mathrm{d}c,\qquad(x^{*}\in\mathcal{X})$ (65)
@@
[…]
@@
$\begin{split}&\pi(x_{\ddot{\tau}}|x_{\dot{\tau}},y_{\dot{\tau}},\theta)\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(x_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(x_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau})})\\ &\quad=\operatorname{\mathrm{N}}\left[x_{\ddot{\tau}};x_{\dot{\tau}},(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(x_{\dot{\tau}},y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau}]}).\end{split}$ (71)
@@
and changing variables back to $V_{\ddot{\tau}}$, the expression is
@@
$\qquad\times\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}(x_{(\dot{\tau},\ddot{\tau})},\eta_{\theta}(v_{\ddot{\tau}})).$
@@
$\zeta_{\theta}(x_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\left\{x_{t}-\eta_{\theta}(v_{\dot{\tau}})-(\eta_{\theta}(v_{\ddot{\tau}})-\eta_{\theta}(v_{\dot{\tau}}))\frac{t-\dot{\tau}}{\ddot{\tau}-\dot{\tau}}\right\}/\rho_{\theta}(y_{\dot{\tau}}),\quad(t\in(\dot{\tau},\ddot{\tau}))$ (74)
@@
$\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\eta_{\theta}(v_{\dot{\tau}})+(\eta_{\theta}(v_{\ddot{\tau}})-\eta_{\theta}(v_{\dot{\tau}}))\frac{t-\dot{\tau}}{\ddot{\tau}-\dot{\tau}}+\rho_{\theta}(y_{\dot{\tau}})z_{t}.\quad(t\in(\dot{\tau},\ddot{\tau}))$ (75)
@@
$\begin{split}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)&=|\eta_{\theta}^{\prime}(v_{\ddot{\tau}})|\operatorname{\mathrm{N}}\left[\eta_{\theta}(v_{\ddot{\tau}});\eta_{\theta}(v_{\dot{\tau}}),(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]\\ &\qquad\times\frac{\mathop{}\!\mathrm{d}\mathbb{X}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}{\mathop{}\!\mathrm{d}\mathbb{M}|(X_{\dot{\tau}}=\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}},\theta)}(\zeta_{\theta}^{-1}(z_{(\dot{\tau},\ddot{\tau})};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),\eta_{\theta}(v_{\ddot{\tau}}))\\ &=|\eta_{\theta}^{\prime}(v_{\ddot{\tau}})|\operatorname{\mathrm{N}}\left[\eta_{\theta}(v_{\ddot{\tau}});\eta_{\theta}(v_{\dot{\tau}}),(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]\\ &\qquad\times\exp\left\{\rho_{\theta}^{-2}(y_{\dot{\tau}})(\Delta_{\theta}(\eta_{\theta}(v_{\ddot{\tau}}),y_{\dot{\tau}})-\Delta_{\theta}(\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}}))\right\}\\ &\qquad\times\exp\left\{-\int_{\dot{\tau}}^{\ddot{\tau}}\varphi_{\theta}(\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),y_{\dot{\tau}})\mathop{}\!\mathrm{d}t\right\},\end{split}$ (77)
@@ Appendix B Localizing and Adapting the Hidden Data Update
[…]
@@
[…]
@@
[…]
@@
merely involve the backward and forward law $\pi(y_{[0,\nu_{1})}^{\dagger}|y_{\nu_{1}},\lambda)$ and $\pi(y_{(\nu_{|\nu|},\omega]}^{\dagger}|y_{\nu_{|\nu|}},\lambda)$ respectively. Simulation of the Markov jump process bridges from $\pi(y_{(\dot{\nu},\ddot{\nu})}^{\dagger}|y_{\{\dot{\nu},\ddot{\nu}\}},\lambda)$ may be carried out according to any of the schemes proposed in [26]. We further observe the decomposition
@@
$\begin{split}&\frac{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\},\{v_{r},z,y\})}{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\{v_{r},z,y\},\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\})}\\ &\quad=\prod_{(\dot{\varsigma}\sim\ddot{\varsigma})\in\varsigma\cap\{0,\omega\}}\frac{\kappa(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}|v_{s},y_{\varsigma})}{\kappa(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger}|v_{s},y_{\varsigma})}\frac{\pi(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger}|v_{s},y_{\varsigma},\theta,\lambda)}{\pi(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger}|v_{s},y_{\varsigma},\theta,\lambda)}\\ &\quad=\prod_{(\dot{\varsigma}\sim\ddot{\varsigma})\in\varsigma\cap\{0,\omega\}}\frac{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger},\gamma_{(\dot{\varsigma},\ddot{\varsigma})})}{\alpha_{(V_{R},Z,Y|y_{\varsigma})}(\gamma_{(\dot{\varsigma},\ddot{\varsigma})},\gamma_{(\dot{\varsigma},\ddot{\varsigma})}^{\dagger})},\end{split}$ (85)